Consider Iran and Iraq and their production of oil and olive oil. Relatively recent OPEC estimates indicate that in July 2012, Iran produced about 4.1 million barrels of oil per day and Iraq produced about 3.2 million barrels of oil per day, making them the second- and third-largest oil producers in OPEC, behind Saudi Arabia (and the 4th and 7th largest oil producing countries in the world).  Suppose that Iran and Iraq both produce barrels of oil and bottles of olive oil, which are sold for the same price in both countries. These are the combinations of the two goods that each country can produce in one day using the same amounts of capital and labor (with all measures in millions throughout):

Iran

Iraq

Oil (Barrels)

Olive Oil (Bottles)

Oil (Barrels)

Olive Oil (Bottles)

0

12

0

8

2

9

2

6

4

6

4

4

6

3

6

2

8

0

8

0

Suppose that without specialization, Iran produces 4 barrels of oil and 6 bottles of olive oil, and Iraq produces 4 barrels of oil and 4 bottles of olive oil (in millions per day).  Are the countries better off or worse off with specialization and trade? Suppose that the terms of trade with specialization are 4 barrels of oil for 5.0 bottles of olive oil, and that 4 barrels of oil are indeed traded for 5.0 bottles of olive oil.  With specialization and trade, how many additional barrels of olive oil can Iraq consume?   

Provide your answer as a number measured in millions rounded to two decimal places. Use "-" to indicate negative amounts. Do not include any symbols, such as "$," "=," "%," or "," in your answer.
Answer #1

Iran: opportunity cost of producing 1 barrel of oil = 12/8 = 1.5 bottles of olive oil

Iraq: opportunity cost of producing 1 barrel of oil = 8/8 = 1 bottle of olive oil

Therefore, Iraq has a comparative advantage in producing of oil (lower opportunity cost)

Before trade:

Total production of oil = 4 + 4 = 8 mn barrels

Total production of olive oil = 4 + 6 = 10 mn bottles

After specialization:

Total production of oil = 8 mn barrels (produced by Iraq only)

Total production of olive oil = 12 mn bottles (produced by Iran only)

Therefore, the countries can be better off with specialization and trade (more production of olive oil).

After trade:

Iraq would consume = 5 - 4 = 1 mn bottles of olive oil. (4 mn barrels of oil are exchanged for 5 mn bottles of olive oil)